# Energy-Efficient Designs for SIM-Based Broadcast MIMO Systems

### v1

@@ II-B Dirty Paper Coding (DPC)
In a multi-user broadcast system, the received signal at user $k$ is given by
@@
where $\mathbf{H}_{k}\in\mathbb{C}^{N_{r}\times N_{t}}$ is the channel matrix for user $k$, $\mathbf{s}_{k}\in\mathbb{C}^{N_{t}\times 1}$ is the transmitted signal intended for user $k$, and $\mathbf{s}_{j}\in\mathbb{C}^{N_{t}\times 1}$ for $j\neq i$ are the transmitted signals intended for the other users, which act as interference for the detection of $\mathbf{s}_{k}$. The noise vector $\mathbf{n}_{k}\in\mathbb{C}^{N_{r}\times 1}$ consists of independent and identically distributed (i.i.d.) elements that are distributed according to $\mathcal{CN}(0,\sigma^{2})$, where $\sigma^{2}$ is the noise variance. DPC is capable of eliminating the interference term $\sum_{j<k}\mathbf{H}_{k}\mathbf{x}_{j}$, which is caused by users $11,2,\dots,k-$ 1. Therefore, the achievable rate of user $k$ is given by
@@
where, by slight abuse of notation, $\mathbf{H}_{k}$ stands for $\mathbf{H}_{k}/\sigma$ (i.e., $\mathbf{H}_{k}$ is normalized by the square root of the noise power), $\mathbf{Q}_{k}=\mathbb{E}\bigl{\{}\mathbf{s}_{k}\mathbf{s}_{k}^{H}\bigr{\}}\succeq\mathbf{0}$ is the input covariance matrix of user $k$ and $\mathbf{Q}=(\mathbf{Q}_{1},\mathbf{Q}_{2},\dots,\mathbf{Q}_{K})$. These covariance matrices are constrained by the total power budget as
@@
II-C Multi-user MIMO with Linear Precoding
Although DPC is a capacity achieving scheme, it has high complexity due to its nonlinear processing nature. On the other hand, linear precoding is much simpler to implement in practice. For linear precoding, the transmitted signal is expressed as
@@ II-C Multi-user MIMO with Linear Precoding
$\mathbf{y}_{k}$ $=\mathbf{H}_{k}\mathbf{x}+\mathbf{n}_{k}$
$=\mathbf{H}_{k}\mathbf{P}_{k}\mathbf{s}_{k}+\sum\nolimits_{j=1,j\neq k}^{K}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{s}_{j}+\mathbf{n}_{k}.$ (8)
@@
×$\times\left(\mathbf{I}+\sum\nolimits_{j=1,j\neq k}^{K}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{P}_{j}^{H}\mathbf{H}_{k}^{H}\right)^{-1}\bigg{|}$ (9)
where $\mathbf{H}_{k}=\mathbf{H}_{k}/\sigma$ and the precoding matrices $\mathbf{P}=(\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{K})$ have to satisfy the total power constraint:
@@ II-D Problem Formulation
In this paper, our goal is to maximize the EE of the considered communication system, which is defined as the ratio of the sum-rate and the total power consumption. To this end, we model the total power consumption as
$P_{\mathrm{tot}}=P_{t}+N_{t}P_{c}+P_{0}+LNP_{s},$ (11)
where $P_{t}$ is the data-dependent transmit signal power, $P_{c}$ is the circuit power per RF chain, $P_{0}$ is the basic power consumed at the BS, and $P_{s}$ is the power consumption of the switching circuits (e.g., PIN diode, varactor diodes) of every SIM meta-element. [^2] In this model, we assume that the control and driving circuits of the SIM are integrated in the BS, and the power consumption of these circuits is already included in the BS power consumption, $P_{0}$. Note that $P_{t}=\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{Q}_{k})$ for the DPC-based scheme and $P_{t}=\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{P}_{k}\mathbf{P}_{k}^{H})$ for the linear precoding scheme. Since BSs have the largest power consumption in mobile networks, the users’ consumed power is not taken into account in the considered EE optimization.
@@
$\underset{\mathbf{Q},\boldsymbol{\phi}}{\operatorname{maximize}}$ $\eta_{\mathrm{dpc}}=\frac{W\sum_{k=1}^{K}R_{\text{BC},k}(\mathbf{Q},\boldsymbol{\phi})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{Q}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (12a)
$\operatorname{subject~{}to}$ $\left|\boldsymbol{\phi}\right|=1,$ (12b)
@@
where $W$ is the system bandwidth. Similarly, the EEmax problem with linear precoding is written as
$\underset{\mathbf{P},\boldsymbol{\phi}}{\operatorname{maximize}}$ $\eta_{\mathrm{lp}}=\frac{W\sum_{k=1}^{K}R_{\text{L},k}(\mathbf{P},\boldsymbol{\phi})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{P}_{k}\mathbf{P}_{k}^{H})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (13a)
$\operatorname{subject~{}to}$ $\left|\boldsymbol{\phi}\right|=1,$ (13b)
@@
Since $W$ is constant, we will drop it when solving (12) and (13), but it is included in simulation results in Section VII. In the following sections, we present our proposed methods for solving the above two EEmax problems.
@@ III Proposed Solution to DPC-based SIM
To solve (12), we present an iterative optimization algorithm which optimizes the covariance matrices and the SIM phase shifts in an alternating manner, which is a prevailing method in existing studies for SIM. In particular, for fixed phase shifts, we propose a novel method which can optimize the covariance matrices in parallel, using a closed-form expression. Our proposed method is derived by applying Dinkelbach’s method to maximize a quadratic lower bound of the objective, iteratively. The phase shifts of the meta-elements of the SIM layers are optimized by a gradient-based optimization method, which is a natural choice, considering the extremely large size of the SIM.
@@ III-A Covariance Matrix Optimization
We remark that the objective function in (12a) is neither convex nor concave with respect to the optimization variables. To deal with this, we exploit the well-know duality between BCs and MACs, introduced in [22], which states that the achievable sum-rate of the MIMO BC equals the achievable sum-rate of the dual MIMO MAC. Accordingly, (12) is equivalent to the EEmax problem in the dual MAC, which is expressed as
$\underset{\mathbf{S},\boldsymbol{\phi}}{\operatorname{maximize}}$ $\quad\frac{\ln\left|\mathbf{I}+\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}\right|}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (14a)
$\operatorname{subject~{}to}$ $\quad\left|\boldsymbol{\phi}\right|=1,$ (14b)
$\quad\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k})\leq P;\mathbf{S}_{k}\succeq\mathbf{0},\forall k,$ (14c)
where $\mathbf{H}_{k}^{H}$ is the dual MAC of user $k$. Also, $\mathbf{S}=(\mathbf{S}_{1},\mathbf{S}_{2},\dots,\mathbf{S}_{K})$, where $\mathbf{S}_{k}\in\mathbb{C}^{N_{r}\times N_{r}}$ is the input covariance matrix of user $k$ in the dual MAC. Note that the equality constraint in (14b) is treated element-wise. The key idea to develop efficient solutions to (12a) is first to drop the power constraint (12c), which results in
[…]
To appreciate the novelty of our proposed method, we briefly describe the block-coordinate method proposed in [23], which optimizes each $\mathbf{S}_{k}$ sequentially, while other variables are fixed. More precisely, let $\mathbf{S}^{(n)}=(\mathbf{S}_{1}^{(n)},\mathbf{S}_{2}^{(n)},\dots,\mathbf{S}_{K}^{(n)})$ denote the current iterate. Then the next iterate $\mathbf{S}_{k}^{(n+1)}$ is obtained as
@@
where $\lambda_{D}$ is a non-negative parameter. For a given $\lambda_{D}$, the above problem can be solved in closed-form [23]. However, such a method requires to find the inverse of $\mathbf{Z}_{k}\in\mathcal{C}^{N_{t}\times N_{t}}$ which has a complexity of $\mathcal{O}(N_{t}^{3})$ in general, and compute the singular value decomposition (SVD) of $\mathbf{T}_{k}$ which has a complexity of $\mathcal{O}(N_{t}^{2}N_{r})$. Thus, the overall complexity of the method presented in [23] is very high.
In this paper, we propose a more efficient method based on the fact that a stationary solution to (15) is also globally optimal since (15) is a concave-convex program. This motivates us to adopt the SCA framework, which is normally used to find a stationary solution for nonconvex programs. To proceed, since $\mathbf{S}_{k}\succeq\mathbf{0}$, we can write $\mathbf{S}_{k}=\mathbf{U}_{k}^{H}\mathbf{U}_{k}$ where $\mathbf{U}_{k}\in\mathbb{C}^{N_{r}\times N_{r}}$. Thus, (15) is equivalent to the following unconstrained optimization problem:
[…]
where $h(\mathbf{U)}=\ln\left|\mathbf{I}+\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\right|$. It is easy to see that $h(\mathbf{U)}$ can be equivalently rewritten as
$h(\mathbf{U)}=\sum_{j=1}^{K}\ln\frac{\left|\mathbf{I}+\sum_{k=j}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\right|}{\left|\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\right|}$ (19)
$=\sum_{j=1}^{K}\ln\left|\mathbf{I}+\Bigl{(}\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\Bigr{)}^{-1}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\right|$ (20)
$=\sum_{j=1}^{K}\ln\left|\mathbf{I}+\mathbf{U}_{j}\mathbf{H}_{j}\Bigl{(}\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\Bigr{)}^{-1}\mathbf{H}_{j}^{H}\mathbf{U}_{j}^{H}\right|.$ (21)
@@
where $\mathbf{A}=\hat{\mathbf{Y}}^{-1}-(\hat{\mathbf{V}}^{H}\hat{\mathbf{V}}+\hat{\mathbf{Y}})^{-1}$. We remark that the above inequality holds for arbitrary $\mathbf{V}$, $\hat{\mathbf{V}}$, $\mathbf{Y}\succ\mathbf{0}$ and $\hat{\mathbf{Y}}\succ\mathbf{0}$, whose sizes are compatible and the equality occurs when $\mathbf{V}=\hat{\mathbf{V}}$ and $\mathbf{Y}=\hat{\mathbf{Y}}$. In light of the SCA framework, we denote by $\mathbf{U}_{j}^{(n)}$ the value of $\mathbf{U}_{j}$ after $n$ iterations. Now let
Input: $\mathbf{H}$, $\mathbf{S}$, $\lambda_{D}^{(0)}\geq 0$, $m\leftarrow 0$
@@ 4     $n\leftarrow 0$
6     repeat
@@
8          Calculate all $\hat{\mathbf{V}}_{j}^{(n)}$, $\hat{\mathbf{Y}}_{j}^{(n)}$, $\mathbf{A}_{j}^{(n)}$, $\mathbf{B}_{j}^{(n)}$

### v2

@@ II-B Dirty Paper Coding (DPC)
To evaluate an upper bound on achievable EE, we first consider DPC, which achieves the capacity region of a Gaussian MIMO BC. The received signal at user $k$ is given by
@@
where $\mathbf{H}_{k}\in\mathbb{C}^{N_{r}\times N_{t}}$ is the channel matrix for user $k$, $\mathbf{s}_{k}\in\mathbb{C}^{N_{t}\times 1}$ is the transmitted signal for user $k$. The noise vector $\mathbf{n}_{k}\in\mathbb{C}^{N_{r}\times 1}$ consists of independent and identically distributed (i.i.d.) elements that are distributed according to $\mathcal{CN}(0,\sigma^{2})$, where $\sigma^{2}$ is the noise variance.
The BS designs $\mathbf{s}_{k}$ with full (non-causal) knowledge of the interference term $\sum_{j<k}\mathbf{H}_{k}\mathbf{s}_{j}$, which results from transmissions to users $1,2,\dots,k-1$. Therefore, DPC can eliminate this interference from previously encoded users, and thus, only the interference caused by users $k+1,\dots,K$ remains. Based on this, the achievable rate of user $k$ is given by
@@
where, by slight abuse of notation, $\mathbf{H}_{k}$ stands for $\mathbf{H}_{k}/\sigma$ (i.e., $\mathbf{H}_{k}$ is normalized by the square root of the noise power), $\mathbf{Q}_{k}=\mathbb{E}\bigl\{\mathbf{s}_{k}\mathbf{s}_{k}^{H}\bigr\}\succeq\mathbf{0}$ is the input covariance matrix of user $k$, $\mathbf{Q}=[\mathbf{Q}_{1},\mathbf{Q}_{2},\dots,\mathbf{Q}_{K}]$ and $\boldsymbol{\phi}=[(\boldsymbol{\phi}^{1})^{T},(\boldsymbol{\phi}^{2})^{T},\dots,(\boldsymbol{\phi}^{L})^{T}]^{T}$. These covariance matrices are constrained by the total power budget as
@@
II-C Multi-user MIMO with LP
Although DPC is a capacity achieving scheme, it has high complexity due to its nonlinear coding nature. On the other hand, LP is much simpler to implement in practice. For LP, the transmitted signal is expressed as
@@ II-C Multi-user MIMO with LP
$\mathbf{y}_{k}=\mathbf{H}_{k}\mathbf{x}+\mathbf{n}_{k}=\mathbf{H}_{k}\mathbf{P}_{k}\mathbf{s}_{k}+\sum_{j=1,j\neq k}^{K}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{s}_{j}+\mathbf{n}_{k}.$ (7)
@@
$\times\left(\mathbf{I}+\sum\nolimits_{j=1,j\neq k}^{K}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{P}_{j}^{H}\mathbf{H}_{k}^{H}\right)^{-1}\bigg|$ (8)
where $\mathbf{H}_{k}=\mathbf{H}_{k}/\sigma$ and the precoding matrices $\mathbf{P}=[\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{K}]$ have to satisfy the total power constraint:
@@
Note that the achievable rate of user $k$ can alternatively be expressed using covariance matrices by defining $\mathbf{Q}_{k}^{\prime}=\mathbf{P}_{k}\mathbf{P}_{k}^{H}$. However, $\mathbf{Q}_{k}^{\prime}$ differs from $\mathbf{Q}_{k}$ for DPC because of differences in the rate expressions. Some existing solutions impose $\operatorname{rank}(\mathbf{Q}_{k})=N_{r}$ to derive $\mathbf{P}_{k}$, but we do not adopt this approach because optimizing $\{\mathbf{Q}_{k}^{\prime}\}$ introduces additional computational complexity compared to directly optimizing $\{\mathbf{P}_{k}\}$.
@@ II-D Problem Formulation
The objective is to maximize the EE of the considered communication system, which is defined as the ratio of the sum-rate and the total power consumption. The total power consumption is modeled as $P_{\mathrm{tot}}=P_{t}+N_{t}P_{c}+P_{0}+LNP_{s},$ where
$P_{t}=\begin{cases}\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{Q}_{k})&\text{DPC}\\ \sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{P}_{k}\mathbf{P}_{k}^{H})&\text{LP}\end{cases}$ (10)
[…]
@@
$\underset{\mathbf{Q},\boldsymbol{\phi}}{\max}$ $\ \eta_{\mathrm{dpc}}=\frac{W\sum_{k=1}^{K}R_{\text{BC},k}(\mathbf{Q},\boldsymbol{\phi})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{Q}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (11a)
$\mathrm{s.t.}$ $\ \left|\phi_{n}^{l}\right|=1,\forall l,n$ (11b)
@@
where $W$ is the system bandwidth. Similarly, the EEmax problem with LP is written as
$\underset{\mathbf{P},\boldsymbol{\phi}}{\max}$ $\ \eta_{\mathrm{lp}}=\frac{W\sum_{k=1}^{K}R_{\text{L},k}(\mathbf{P},\boldsymbol{\phi})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{P}_{k}\mathbf{P}_{k}^{H})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (12a)
$\mathrm{s.t.}$ $\ \left|\phi_{n}^{l}\right|=1,\forall l,n$ (12b)
@@
Note that the equality constraints in (11b) and (12b) are treated element-wise. Since $W$ is constant, it is dropped when solving (11) and (12), but included in simulation results in Section VII. Next, we present our proposed methods for solving the above two EEmax problems.
@@ III Proposed Solution to DPC-based SIM
To solve (11), we present an iterative optimization algorithm that alternates between optimizing the covariance matrices and the SIM phase shifts, a prevailing method in SIM-related studies. In particular, for fixed phase shifts, we propose a novel method to optimize the covariance matrices in parallel, using a closed-form water-filling solution, which allocates more transmit power to “stronger” user channels. This nice result is achieved by applying Dinkelbach’s method to iteratively maximize a quadratic lower bound of the objective function. The phase shifts of the SIM meta-elements are optimized by a gradient-based optimization method, which is a viable choice, considering the extremely large size of the SIM.
@@ III-A Covariance Matrix Optimization
For ease of readability, we briefly outlined the steps for optimizing covariance matrices as follows:
  • Reformulate (11) using the BCs-MACs duality to obtain a more tractable equivalent formulation.
  • Derive a quadratic lower bound for the achievable sum-rate under the SCA framework.
  • Apply Dinkelbach’s method to solve the fractional EE objective, leading to a water-filling solution for the covariance matrices.
III-A1 Problem Reformulation
The EEmax problem in (11a) is challenging to solve since its objective function is neither convex nor concave with respect to the optimization variables. To deal with this, we exploit the well-known duality between BCs and MACs [21], which states that the capacity region of the MIMO BC is equal to the capacity achieving region of the dual MIMO MAC. For single-user MIMO systems, this duality means that the capacity of a downlink channel $\mathbf{H}$ equals that of a dual uplink channel $\mathbf{H}^{H}$ under the same power constraint, which is easily understood. Extending this idea to multi-user MIMO systems, for a given channel between the BS and user $k$, the BC channel $\mathbf{H}_{k}$ corresponds to the dual MAC channel $\mathbf{H}_{k}^{H}$. Similarly, each transmit covariance matrix $\mathbf{Q}_{k}$ in the BC has as its counterpart a corresponding dual covariance matrix $\mathbf{S}_{k}$ in the dual MAC. [^3] It should be noted that $\mathbf{Q}_{k}$ can not be directly mapped to $\mathbf{S}_{k}$. Instead, it is required to compute first $\mathbf{S}_{K},\mathbf{S}_{K-1},\dots,\mathbf{S}_{k+1}$ before obtaining $\mathbf{S}_{k}$ [21]. Accordingly, (11) is equivalent to the EEmax problem in the dual MAC, which is expressed as
$\underset{\mathbf{S}}{\operatorname{maximize}}$ $\quad\frac{\ln\left|\mathbf{I}+\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}\right|}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (13a)
$\quad\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k})\leq P_{\text{max}};\mathbf{S}_{k}\succeq\mathbf{0},\forall k,$ (13b)
where we have omitted (11b) as the phase shifts are fixed, and the power constraint (13b) has the same maximum power as (11c). To solve (13) efficiently, we temporarily drop the power constraint (11c), which results in
$\!\!\!\text{$\underset{\mathbf{S}_{k}\succeq 0}{\operatorname{maximize}}$ }g(\mathbf{S})=$ $\frac{\ln\left|\mathbf{I}+\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}\right|}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}.$ (14)
To appreciate the novelty of our proposed method, we briefly describe the block-coordinate method proposed in [19], which optimizes each $\mathbf{S}_{k}$ sequentially, while other variables are fixed. More precisely, let $\mathbf{S}^{(n)}=[\mathbf{S}_{1}^{(n)},\mathbf{S}_{2}^{(n)},\dots,\mathbf{S}_{K}^{(n)}]$ denote the current iterate. Then the next iterate $\mathbf{S}_{k}^{(n+1)}$ is obtained as
@@ III-A1 Problem Reformulation
where $\lambda_{D}$ is a non-negative parameter. For a given $\lambda_{D}$, the above problem can be solved in closed-form [19]. However, such a method requires finding the inverse of $\mathbf{Z}_{k}\in\mathcal{C}^{N_{t}\times N_{t}}$ which has a complexity of $\mathcal{O}(N_{t}^{3})$ in general, and computing the singular value decomposition (SVD) of $\mathbf{T}_{k}$ which has a complexity of $\mathcal{O}(N_{t}^{2}N_{r})$. Thus, the overall complexity of the method presented in [19] is very high.
III-A2 SCA Method: A Tight Lower Bound of Sum Rate
We now propose a more efficient method by leveraging the fact that a stationary solution to (14) is also globally optimal since (14) is a concave-convex program. This observation motivates us to adopt the SCA framework, which is typically used to find a stationary solution for nonconvex programs. To proceed, since $\mathbf{S}_{k}\succeq\mathbf{0}$, we can write $\mathbf{S}_{k}=\mathbf{U}_{k}^{H}\mathbf{U}_{k}$ where $\mathbf{U}_{k}\in\mathbb{C}^{N_{r}\times N_{r}}$. Thus, (14) is equivalent to the following unconstrained optimization problem:
$\text{$\underset{\mathbf{U}_{k}}{\operatorname{maximize}}$ }g(\mathbf{U})=\frac{h(\mathbf{U)}}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{U}_{k}^{H}\mathbf{U}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}.$ (17)
where $h(\mathbf{U)}=\ln\left|\mathbf{I}+\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\right|$ and $\mathbf{U}=[\mathbf{U}_{1},\mathbf{U}_{2},\dots,\mathbf{U}_{K}]$. It is easy to see that $h(\mathbf{U)}$ can be equivalently rewritten as
$h(\mathbf{U)}=\sum_{j=1}^{K}\ln\Bigl|\mathbf{I}+\Bigl(\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\Bigr)^{-1}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\Bigr|$
$=\sum_{j=1}^{K}\ln\Bigl|\mathbf{I}+\mathbf{U}_{j}\mathbf{H}_{j}\Bigl(\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\Bigr)^{-1}\mathbf{H}_{j}^{H}\mathbf{U}_{j}^{H}\Bigr|.$
@@ III-A2 SCA Method: A Tight Lower Bound of Sum Rate
where $\mathbf{A}=\hat{\mathbf{Y}}^{-1}-(\hat{\mathbf{V}}^{H}\hat{\mathbf{V}}+\hat{\mathbf{Y}})^{-1}$. We remark that the above inequality holds for arbitrary $\mathbf{V}$, $\hat{\mathbf{V}}$, $\mathbf{Y}\succ\mathbf{0}$ and $\hat{\mathbf{Y}}\succ\mathbf{0}$, provided that their sizes are compatible and the equality occurs when $\mathbf{V}=\hat{\mathbf{V}}$ and $\mathbf{Y}=\hat{\mathbf{Y}}$. In light of the SCA framework, we denote by $\mathbf{U}_{j}^{(n)}$ the value of $\mathbf{U}_{j}$ after $n$ iterations. Let
Input: $\mathbf{H}$, $\mathbf{S}$, $\lambda^{(0)}\geq 0$, $n\leftarrow 0$
@@
6    Calculate $\hat{\mathbf{V}}_{j}^{(n)}$, $\hat{\mathbf{Y}}_{j}^{(n)}$, $\mathbf{A}_{j}^{(n)}$, $\mathbf{B}_{j}^{(n)}$
@@
8    Decompose $\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)}\mathbf{H}_{j}^{H}=\mathbf{P}_{j}\boldsymbol{\Sigma}_{j}\mathbf{P}_{j}^{H}$
@@
12    Solve for $\lambda_{D}$ in (30)
13
14    Compute $\mathbf{U}_{j}^{(n+1)}$ using (27) for $j=1,2,\cdots,K$
15
16    $\lambda^{(n+1)}\leftarrow\lambda_{D}$